# Learning out-of-time-ordered correlators with classical kernel methods

### v1

@@ IV Results
We now report the main numerical results of this study. In Figures 713 we show the root mean-squared-error (RMSE), mean absolute error (MAE) and coefficient of determination ($R^{2}$) performance metrics for all 6 of the kernels considered and all problem instances. The performance metrics shown on the plots are those for the models with the best hyperparameter values found in the cross-validation (see Appendix B). We show in separate plots the metrics evaluated on training and testing datasets. Tables containing all numerical values plotted in Figures 713 (including those not visible on the plots) can be found in Appendix C.
Figure 6: ($\mathbf{\mathscr{O}_{XZ}}$ with $\mathbf{\mathscr{H}_{1}}$) Learning performance metrics for the best models found in the cross-validation. The metrics are calculated based on predictions made by the models on the testing (Test) and training (Train) sets, with labels determined by $\mathscr{O}_{XZ}$ and $\mathscr{H}_{1}$, for 5-40 qubit systems. The RMSE, MAE and $R^{2}$ values are displayed for the testing set in (a)–(c) and the training set in (d)–(f) respectively.
Figure 7: ($\mathbf{\mathscr{O}_{Sum}}$ with $\mathbf{\mathscr{H}_{1}}$) Learning performance metrics for the best models found in the cross-validation. The metrics are calculated based on predictions made by the models on the testing (Test) and training (Train) sets, with labels determined by $\mathscr{O}_{Sum}$ and $\mathscr{H}_{1}$, for 5-40 qubit systems. The RMSE, MAE and $R^{2}$ values are displayed for the testing set in (a)–(c) and the training set in (d)–(f) respectively.
Figure 8: ($\mathbf{\mathscr{O}_{XZ}}$ with $\mathbf{\mathscr{H}_{2}}$) Learning performance metrics for the best models found in the cross-validation. The metrics are calculated based on predictions made by the models on the testing (Test) and training (Train) sets, with labels determined by $\mathscr{O}_{XZ}$ and $\mathscr{H}_{2}$, for 5-40 qubit systems. The RMSE, MAE and $R^{2}$ values are displayed for the testing set in (a)–(c) and the training set in (d)–(f) respectively.
Figure 9: ($\mathbf{\mathscr{O}_{Sum}}$ with $\mathbf{\mathscr{H}_{2}}$) Learning performance metrics for the best models found in the cross-validation. The metrics are calculated based on predictions made by the models on the testing (Test) and training (Train) sets, with labels determined by $\mathscr{O}_{Sum}$ and $\mathscr{H}_{2}$, for 5-40 qubit systems. The RMSE, MAE and $R^{2}$ values are displayed for the testing set in (a)–(c) and the training set in (d)–(f) respectively.
Figure 10: ($\mathbf{\mathscr{O}_{XZ}}$ with $\mathbf{\mathscr{H}_{3}}$) Learning performance metrics for the best models found in the cross-validation. The metrics are calculated based on predictions made by the models on the testing (Test) and training (Train) sets, with labels determined by $\mathscr{O}_{XZ}$ and $\mathscr{H}_{3}$, for 5-40 qubit systems. The RMSE, MAE and $R^{2}$ values are displayed for the testing set in (a)–(c) and the training set in (d)–(f) respectively.
Figure 11: ($\mathbf{\mathscr{O}_{Sum}}$ with $\mathbf{\mathscr{H}_{3}}$) Learning performance metrics for the best models found in the cross-validation. The metrics are calculated based on predictions made by the models on the testing (Test) and training (Train) sets, with labels determined by $\mathscr{O}_{Sum}$ and $\mathscr{H}_{3}$, for 5-40 qubit systems. The RMSE, MAE and $R^{2}$ values are displayed for the testing set in (a)–(c) and the training set in (d)–(f) respectively.
Figure 12: ($\mathbf{\mathscr{O}_{XZ}}$ with $\mathbf{\mathscr{H}_{4}}$) Learning performance metrics for the best models found in the cross-validation. The metrics are calculated based on predictions made by the models on the testing (Test) and training (Train) sets, with labels determined by $\mathscr{O}_{XZ}$ and $\mathscr{H}_{4}$, for 5-40 qubit systems. The RMSE, MAE and $R^{2}$ values are displayed for the testing set in (a)–(c) and the training set in (d)–(f) respectively.
Figure 13: ($\mathbf{\mathscr{O}_{Sum}}$ with $\mathbf{\mathscr{H}_{4}}$) Learning performance metrics for the best models found in the cross-validation. The metrics are calculated based on predictions made by the models on the testing (Test) and training (Train) sets, with labels determined by $\mathscr{O}_{Sum}$ and $\mathscr{H}_{4}$, for the 5-40 qubit systems. The RMSE, MAE and $R^{2}$ values are displayed for the testing set in (a)–(c) and the training set in (d)–(f) respectively.
In all Figures 713, the linear and cosine kernels are not visible on any of the plots, since they perform so poorly that they never achieve performance metrics within the chosen plot ranges. In contrast, the polynomial and sigmoid kernels perform reasonably well. For a handful of the problem instances, the polynomial and sigmoid kernels achieve $R^{2}$ scores on the testing sets exceeding 0.8, and sometimes 0.9, with similar RMSE and MAE values to the Laplacian and RBF kernels. However, all of these kernels are consistently outperformed by the Laplacian and RBF kernels. One can easily see that the performance metrics for the Laplacian and RBF kernels are better on all problem instances considered. For this reason we focus the rest of this section on reporting the performance of the Laplacian and RBF kernels, with an emphasis on the testing performance.
In Figures 7 and 7, we can see the performance of the Laplacian and RBF kernels (together with the other kernels) on both the testing and training sets for $\mathscr{O}_{XZ}$ and $\mathscr{O}_{Sum}$, respectively, with $\mathscr{H}_{1}$. For $\mathscr{O}_{XZ}$ with $\mathscr{H}_{1}$, the Laplacian kernel on the testing set achieves RMSE values between 0.0625 and 0.1137, MAE values from 0.0388 to 0.0763, and $R^{2}$ scores ranging from 0.9228 to 0.9696. The RBF kernel, in comparison, shows RMSE values from 0.0837 to 0.1359, MAE values between 0.0431 and 0.0790, and $R^{2}$ scores ranging from 0.8771 to 0.9646. For $\mathscr{O}_{Sum}$ with $\mathscr{H}_{1}$, the Laplacian kernel on the testing set yields RMSE values from 0.6176 to 0.8534, MAE values between 0.3579 and 0.5653, and $R^{2}$ scores from 0.9429 to 0.9652. While the RBF kernel shows RMSE values from 0.7402 to 1.1179, MAE values between 0.4385 and 0.6916, and $R^{2}$ scores ranging from 0.8979 to 0.9556. In both cases, the training performance for both kernels is consistently better than the testing performance, with the Laplacian kernel approaching near-perfect training performance.
Figures 9 and 9 present similar results for $\mathscr{O}_{XZ}$ and $\mathscr{O}_{Sum}$, respectively, with $\mathscr{H}_{2}$. For $\mathscr{O}_{XZ}$ with $\mathscr{H}_{2}$, the Laplacian kernel on the testing set achieves RMSE values between 0.1100 and 0.1372, MAE values from 0.0709 to 0.0920, and $R^{2}$ scores ranging from 0.8452 to 0.8982. Comparatively, the RBF kernel shows RMSE values between 0.0782 and 0.1399, MAE values from 0.0457 to 0.0894, and $R^{2}$ scores ranging from 0.8306 to 0.9656. For $\mathscr{O}_{Sum}$ with $\mathscr{H}_{2}$, the Laplacian kernel on the testing set yields RMSE values from 0.8730 to 1.0664, MAE values between 0.5915 and 0.7712, and $R^{2}$ scores from 0.8862 to 0.9108. Similarly, the RBF kernel shows RMSE values between 0.6089 and 0.9708, MAE values from 0.3170 to 0.6700, and $R^{2}$ scores ranging from 0.8957 to 0.9706. As before, the training performance for both kernels is always superior to the testing performance, but is marginally worse compared with Figures 7 and 7.
Figures 11 and 11 provide results for $\mathscr{O}_{XZ}$ and $\mathscr{O}_{Sum}$ with $\mathscr{H}_{3}$. For $\mathscr{O}_{XZ}$ with $\mathscr{H}_{3}$, the Laplacian kernel on the testing set achieves RMSE values from 0.0670 to 0.0843, MAE values between 0.0424 and 0.0573, and $R^{2}$ scores from 0.9537 to 0.9745. The RBF kernel shows comparable RMSE values ranging from 0.0459 to 0.0724, MAE values from 0.0290 to 0.0511, and $R^{2}$ scores from 0.9682 to 0.9866. For $\mathscr{O}_{Sum}$ with $\mathscr{H}_{3}$, the Laplacian kernel on the testing set yields RMSE values between 0.5740 and 0.6719, MAE values from 0.3564 to 0.4879, and $R^{2}$ scores ranging from 0.9606 to 0.9764. Meanwhile the RBF kernel shows RMSE values between 0.3831 and 0.6243, MAE values from 0.2247 to 0.4376, and $R^{2}$ scores ranging from 0.9703 to 0.9882. As with previous cases, the training performance for both kernels exceeds their testing performance. However, in these instances the Laplacian kernel exhibits perfect training performance to at least 4 decimal places.
Finally, Figures 13 and 13 show the performance of the kernels on the datasets for $\mathscr{O}_{XZ}$ and $\mathscr{O}_{Sum}$ with $\mathscr{H}_{4}$. For $\mathscr{O}_{XZ}$ with $\mathscr{H}_{4}$, the Laplacian kernel on the testing set achieves RMSE values from 0.0934 to 0.1282, MAE values between 0.0514 and 0.0715, and $R^{2}$ scores ranging from 0.7511 to 0.9214. The RBF kernel shows RMSE values from 0.0888 to 0.1499, MAE values between 0.0468 and 0.0731, and $R^{2}$ scores from 0.7167 to 0.8931. For $\mathscr{O}_{Sum}$ with $\mathscr{H}_{4}$, the Laplacian kernel yields RMSE values from 0.5859 to 0.9806, MAE values between 0.3555 and 0.5750, and $R^{2}$ scores ranging from 0.8103 to 0.9607. While the RBF kernel shows RMSE values from 0.7614 to 1.0930, MAE values between 0.3943 and 0.5934, and $R^{2}$ scores from 0.7908 to 0.9091. As in all the other cases, the training performance for both kernels is consistently better than their testing performance, with the Laplacian kernel again achieving almost perfect training performance.
@@
With the exception of a few cases, namely $n=15,$ 20 and 25 in Figure 13, and $n=10$ and $25$ in Figure 13, the $R^{2}$ scores on the testing data always exceed 0.8 for both of the Laplacian and RBF kernels, and often exceed 0.9. Similarly, the RMSE and MAE for the Laplacian and RBF kernels on the testing data are small. In the problem instances associated with $\mathscr{O}_{XZ}$, they almost never exceed 0.15 and are often less than 0.1. In the problem instances associated with $\mathscr{O}_{Sum}$, they almost never exceed 1.5 and are often less than 1.0. This indicates that the numerical error in the labels predicted by at least one of these kernels for $\mathscr{O}_{XZ}$, and $\mathscr{O}_{Sum}$, will generally be smaller than 5–$7.5\%$, and 4–$8\%$, of the range of the functions which returns values in $[-1,1]$, and $[-3,9]$, respectively. Note though, that the smallest numerical value of any label determined by $\mathscr{O}_{XZ}$, and $\mathscr{O}_{Sum}$, present in the datasets was -0.2297, and -0.1086, respectively.
@@
Overall, the results show that the learning performance metrics for the RBF and Laplacian are similar and consistently the best for all problem instances considered. However, while the Laplacian kernel seems to perform better than the RBF kernel in many instances, it also exhibits a larger difference between training and testing performance. This could be interpreted as partial overfitting, which is also supported by observing that the best regularisation strength among those trialled for the Laplacian kernel was frequently $\lambda=0$ (see Appendix B).

### v2

@@ IV Results
Figure 4: Coefficient of determination ($R^{2}$) for the models trained on all 1000 training datapoints with the best hyperparameter values trialled in the cross-validation, making predictions on the testing sets. The top row of plots shows the results for the datasets with labels determined by $\mathscr{O}_{XZ}$ with (a) $\mathscr{H}_{1}$, (b) $\mathscr{H}_{2}$, (c) $\mathscr{H}_{3}$ and (d) $\mathscr{H}_{4}$. The bottom row of plots shows the results for the datasets with labels determined by $\mathscr{O}_{Sum}$ with (e) $\mathscr{H}_{1}$, (f) $\mathscr{H}_{2}$, (g) $\mathscr{H}_{3}$ and (h) $\mathscr{H}_{4}$.
We now report the main numerical results of this study. In Figure 4 we plot the coefficient of determination ($R^{2}$) performance metric for the RBF, Laplacian, polynomial, and sigmoid kernels based on predictions made on the testing sets for all problem instances. The linear and cosine kernels are not included in the figure because they perform so poorly that adjusting the plot ranges to include them would obscure details in the important $R^{2}\in[0.8,1]$ range. Table I presents a summary of the $R^{2}$, root-mean-squared-error (RMSE), and mean-absolute-error (MAE) performance metrics obtained with the Laplacian and RBF kernels, which consistently achieved the best performance metrics of all the kernels we consider. Additional numerical results, including the $R^{2}$, RMSE, and MAE values for all six kernels, making predictions on both the training and testing sets, are provided in Tables D.1, D.2, D.3, D.4, D.5, D.6, D.7, and D.8 of Appendix D. All results in Appendix D (which includes the $R^{2}$ values in Figure 4, and the results which are summarised in Table I) are obtained using models trained on all 1000 training datapoints with the best hyperparameter values trialled in the cross-validation (see Section III.2). In Figure 5, we also provide results showing how the kernels perform when supplied less than 1000 training datapoints.
Since the linear and cosine kernels perform poorly, we will not discuss their performance in detail, but instead provide reasons explaining their poor performance. Specifically, we can combine (7) and (26) to show that the linear kernel can only be used to learn models of the form,
$f(x)=\left\langle x,\sum_{i=1}^{M}\alpha_{i}\mathbf{x}_{i}\right\rangle.$ (33)
Similarly, combining (7) and (31) shows that the cosine model can only be used to learn models of the form,
$f(x)=\left\langle\frac{x}{\|x\|},\sum_{i=1}^{M}\alpha_{i}\frac{\mathbf{x}_{i}}{\|\mathbf{x}_{i}\|}\right\rangle.$ (34)
From (33) is is clear that the linear kernel can only be used to express linear functions, while (34) shows that the cosine kernel can only be used to express functions which are independent of the Euclidean norm of their argument. However, from (24) and (25), one can see that $\mathscr{O}_{XZ}$ and $\mathscr{O}_{Sum}$ are highly non-linear and depend substantively on the Euclidean norm of their argument (see Section III.1). For this reason, we should not expect the linear and cosine kernels to perform well on the learning problems considered in this work.
On the other hand, the polynomial and sigmoid kernels perform reasonably well. As can be seen in Figure 4, at least one of these kernels achieves an $R^{2}$ score on the testing sets exceeding 0.8 in 37 of the 64 problem instances, with similar RMSE and MAE values to the Laplacian and RBF kernels (see Appendix D). Note though, that the sigmoid kernel performs poorly at predicting both $\mathscr{O}_{XZ}$ and $\mathscr{O}_{Sum}$ with $\mathscr{H}_{4}$ for system sizes of $n\in\{30,40\}$ (see Figure 4(d) and 4(h)), achieving large negative $R^{2}$ scores which fall outside the plot ranges. Despite the overall good performance of the polynomial and sigmoid kernels, it is clear from Figure 4 that the Laplacian and RBF kernels consistently perform the best. For this reason we focus the rest of this section primarily on reporting the performance of the Laplacian and RBF kernels.
$\boldsymbol{R^{2}}$ RMSE MAE Function Kernel Hamiltonians Min Mean SD Max Mean SD Max Mean SD $\mathscr{O}_{XZ}$ $\mathcal{K}_{\textrm{Lap}}$ $\mathscr{H}_{1}$ 0.9228 0.9440 0.0136 0.1137 0.0930 0.0150 0.0763 0.0562 0.0112 $\mathscr{H}_{2}$ 0.8452 0.8785 0.0188 0.1372 0.1229 0.0105 0.0920 0.0805 0.0075 $\mathscr{H}_{3}$ 0.9537 0.9686 0.0075 0.0843 0.0714 0.0060 0.0573 0.0483 0.0055 $\mathscr{H}_{4}$ 0.7511 0.8298 0.0552 0.1282 0.1085 0.0110 0.0715 0.0603 0.0073 $\mathcal{K}_{\textrm{RBF}}$ $\mathscr{H}_{1}$ 0.8771 0.9142 0.0273 0.1359 0.1132 0.0174 0.0790 0.0634 0.0123 $\mathscr{H}_{2}$ 0.8306 0.8865 0.0372 0.1399 0.1158 0.0173 0.0894 0.0766 0.0132 $\mathscr{H}_{3}$ 0.9682 0.9785 0.0054 0.0724 0.0593 0.0079 0.0511 0.0393 0.0064 $\mathscr{H}_{4}$ 0.7167 0.8112 0.0550 0.1499 0.1164 0.0201 0.0731 0.0585 0.0097 $\mathscr{O}_{Sum}$ $\mathcal{K}_{\textrm{Lap}}$ $\mathscr{H}_{1}$ 0.9429 0.9558 0.0066 0.8534 0.7363 0.0687 0.5653 0.4475 0.0682 $\mathscr{H}_{2}$ 0.8862 0.8985 0.0097 1.0664 0.9749 0.0713 0.7712 0.6841 0.0661 $\mathscr{H}_{3}$ 0.9606 0.9716 0.0058 0.6719 0.5973 0.0302 0.4879 0.3979 0.0402 $\mathscr{H}_{4}$ 0.8103 0.8697 0.0505 0.9806 0.8370 0.1279 0.5750 0.4868 0.0760 $\mathcal{K}_{\textrm{RBF}}$ $\mathscr{H}_{1}$ 0.8979 0.9214 0.0180 1.1179 0.9772 0.1227 0.6916 0.5689 0.0821 $\mathscr{H}_{2}$ 0.8957 0.9109 0.0234 0.9708 0.8939 0.1096 0.6700 0.5925 0.1095 $\mathscr{H}_{3}$ 0.9703 0.9822 0.0055 0.6243 0.4737 0.0820 0.4376 0.3113 0.0693 $\mathscr{H}_{4}$ 0.7908 0.8589 0.0420 1.0930 0.8920 0.1084 0.5934 0.4860 0.0740
Table I: Summary of the $R^{2}$, RMSE, and MAE values for the Laplacian and RBF kernels on the testing sets with labels determined by $\mathscr{O}_{XZ}$ or $\mathscr{O}_{Sum}$, and $\mathscr{H}_{1}$, $\mathscr{H}_{2}$, $\mathscr{H}_{3}$, or $\mathscr{H}_{4}$. The table includes the smallest $R^{2}$, highest RMSE, and highest MAE, together with the mean and standard deviation (SD) of the eight values of each metric obtained across eight system sizes ($n\in\{5,10,15,\ldots,40\}$). For each function, the best value of each quantity reported in the table is written in boldface, with a smaller SD considered better so that the reported means are more statistically representative of expected performance. Note that the models used to obtain the results summarised here are trained on all 1000 training datapoints with the best hyperparameter values trialled in the cross-validation.
In Table I, we provide a summary of the $R^{2}$, RMSE, and MAE values obtained with the Laplacian and RBF kernels on the testing sets with labels determined by $\mathscr{O}_{XZ}$ or $\mathscr{O}_{Sum}$, and one of the parameterised sets of Hamiltonians $\mathscr{H}_{1}$, $\mathscr{H}_{2}$, $\mathscr{H}_{3}$, or $\mathscr{H}_{4}$. Specifically, the table contains the smallest $R^{2}$, highest RMSE, and highest MAE, together with the mean and standard deviation of the eight values of each metric obtained across the eight system sizes $n\in\{5,10,15,\ldots,40\}$. These results offer an overview of how well the Laplacian and RBF kernels perform in learning the two OTOC functions, $\mathscr{O}_{XZ}$ and $\mathscr{O}_{Sum}$, across a range of quantum systems with various system sizes up to 40 qubits. This provides insight into the kernels’ suitability for the learning problems under study.
From Table I, we see that in the problem instances associated with $\mathscr{O}_{XZ}$, the $R^{2}$ scores achieved by the Laplacian and RBF kernels on the testing sets are all at least 0.7167, with averages over the eight system sizes lying between 0.8112 and 0.9785 for the different sets of Hamiltonians. Similarly, for the problem instances associated with $\mathscr{O}_{Sum}$, the $R^{2}$ achieved by the kernels are at least 0.7908, with averages lying between 0.8598 and 0.9822 for the various Hamiltonians. We also see that the collections of $R^{2}$ scores are distributed with small standard deviations, the greatest of which is 0.0552. This indicates that the $R^{2}$ scores are distributed closely around their associated means, which is consistent with Figure 4.
@@
Table I also shows that, for the problem instances associated with $\mathscr{O}_{XZ}$, the RMSE and MAE values all lie below 0.1499 and 0.0920 respectively. While those for the problem instances associated with $\mathscr{O}_{Sum}$ all lie below 1.1179 and 0.7712 respectively. This indicates that the numerical error in the labels predicted by either one of the kernels for $\mathscr{O}_{XZ}$, or $\mathscr{O}_{Sum}$, will on average be much smaller than $7.5\%$, or $9.3\%$, of the range of the functions which return values in $[-1,1]$, or $[-3,9]$, respectively. Note though, that the smallest numerical value of any label determined by $\mathscr{O}_{XZ}$, and $\mathscr{O}_{Sum}$, present in the datasets was -0.2297, and -0.1086, respectively.
Figure 5: Average coefficient of determination (Average $R^{2}$) for each kernel making predictions on the 40-qubit testing sets, plotted against the number of training data samples $M$ used to train the associated ML models. Each point in the plots corresponds to the numerical value of $R^{2}$ averaged over 20 different models using the same kernel. Each of the models is the result of using a random subset of the training data containing $M$ of the total 1000 training data samples. The hyperparameters are fixed to be the best hyperparameters found during the 10-fold cross-validation performed on all 1000 training data samples. Plots (a), (b), (c) and (d) show the average $R^{2}$ score on the testing datasets with labels determined by $\mathscr{O}_{XZ}$ with $\mathscr{H}_{1}$, $\mathscr{H}_{2}$, $\mathscr{H}_{3}$ and $\mathscr{H}_{4}$, respectively. Plots (e), (f), (g) and (h) show the average $R^{2}$ score on the testing datasets with labels determined by $\mathscr{O}_{Sum}$ with $\mathscr{H}_{1}$, $\mathscr{H}_{2}$, $\mathscr{H}_{3}$ and $\mathscr{H}_{4}$, respectively.
Another aspect of the results which warrants mention are the variations in the best trialled values of each hyperparameter (see Tables C.2, C.3, C.4, and C.5 of Appendix C), and the performance of each kernel, as the system size $n$ changes. For the kernels which did not perform well (e.g., the linear and cosine kernels), one would not expect the hyperparameter values or performance metrics to change in a predictable way. However for the Laplacian and RBF kernels, which both depend on two hyperparameters $\lambda$ and $\gamma$, one might expect the optimal choice of these hyperparameters and the associated performance metrics to change in a predictable manner. Unfortunately, this does not appear to be the case, which raises the question of the underlying causes of these variations.
One factor of our method which likely contributes to these variations is the decreasing density of input training data samples with increasing system size $n$. This happens because we sample the input training data from a sphere of radius proportional to $n$, while the total number of training data samples remains fixed at 1000. The decrease in density likely impacts the value of $\gamma$, since $\gamma$ controls the rate of exponential decay of the terms in (7), which in turn affects how the model interpolates between the input training data samples. In order to establish a clearer trend for the hyperparameter values, and hence the associated performance metrics, we expect that maintaining the density of training data samples would assist. However, this would require a number of training data samples scaling cubically with the system size (as does the volume of the sphere from which the data is sampled), which was not feasible with our computational budget.
@@
Similarly, while the Laplacian kernel may appear to perform better than the RBF kernel in many instances, it also exhibits a larger difference between training and testing performance (see Appendix D). This could be interpreted as partial overfitting, which can usually be remedied by using a small positive value of $\lambda$. However the best regularisation strength among those trialled for the Laplacian kernel was frequently $\lambda=0$ (see Appendix C). This, and the absence of a clear trend in the best trialled hyperparameter values, suggests that a more finely grained grid search may improve the results. This would allow hyperparameter values which are closer to optimal to be found, which may help in uncovering the underlying trend that one might expect to observe in the optimal values. However we did not perform this here in order to keep the costs of training manageable (see (36)).
Given the cost of producing training data, it is natural to ask whether one can obtain similar results with fewer data samples. To investigate this, we performed a collection of learning experiments which involve varying the amount of training data supplied to the models. First, we set the hyperparameter values for each kernel in each problem instance equal to the best values trialled in the cross-validation performed on all of the training data (see Appendix C). Next, we choose different training dataset sizes $M\in\mathbb{N}$ to supply to the models. Specifically, for each $M\in\{50m:m=1,2,\ldots,20\}$, we randomly sample $M$ of the 1000 training data points, and repeat this 20 times to obtain 20 (possibly intersecting) subsets of size $M$. We then train a model on each of these 20 subsets of the training data. Finally, we calculate the average $R^{2}$ score achieved by the 20 models making predictions on the entire testing set, as an indicator of how the models perform with only $M$ training data samples. The results of this analysis are shown in Figure 5 for the 40-qubit problem instances. The plots show that, even with only a few hundred training data points, the average $R^{2}$ scores for the Laplacian and RBF kernels often remain above $0.8$, and occasionally reach above $0.9$. This shows that the kernels are capable of obtaining good learning performance metrics when supplied with fewer training data samples than were used in the earlier analysis.
Overall, the results show that the learning performance of the RBF and Laplacian kernels are similar and consistently the best (see Figure 4). On all problem instances, they obtain good values for all three performance metrics considered (see Table I) after tuning their hyperparameters with a standard and inexpensive cross-validation. Even when supplied with just a few hundred training data samples, the RBF and Laplacian kernels manage to achieve $R^{2}$ scores frequently in excess of 0.8 (see Figure 5). We also note that the results indicate that the performance of the Laplacian and RBF kernels depends more on the choice of Hamiltonian than the OTOC function we try to learn. For example, it is clear from Table I that the kernels perform better at learning OTOCs describing $\mathscr{H}_{3}$ than OTOCs describing the other sets.